OpenStudy (anonymous):

The point (3, -4) is on the terminal side of an angle theta . What is cos theta ? A. -3/5 B.3/5 C.-3/4 D.3/4 E. 3

OpenStudy (anonymous):

cos = x/r(or hypo)

OpenStudy (anonymous):

how do you find r?

OpenStudy (anonymous):

r^2=a^2+b^2=c^2(this is r) r^2= 3^2 + (-4)^2 = c^2